The main uptake air duct of a forced air gas heater is 0.4 m in diameter. What is the average speed of air in the duct if it carries a volume equal to that of the house’s interior every 13 min? The inside volume of the house is equivalent to a rectangular solid 10.29 m wide by 11.11 m long by 2.8 m high.


Round your answer to 2 decimal places


1
Expert's answer
2021-04-23T11:06:17-0400

Volumetric flow rate"(Q)", is given as,

"Q=Av"

"A:" Area of cross section of duct

"v:" velocity of flow


Also,

"Q=\dfrac{V}{t}"

"Av=\dfrac{V}{t}"

"v=\dfrac{V}{At}=\dfrac{LBH}{\dfrac{\pi d^2}{4}t}"

"v=\dfrac{10.29\times11.11\times2.8}{\dfrac{\pi\times(0.4)^2}{4}\times(13\times60)}"

"v=3.26\space m/s"